# Insights in $f(Q)$ cosmology: the relevance of the connection

### v2

@@ I Introduction
The standard model of cosmology, known as the $\Lambda$CDM model, has been remarkably successful in describing a wide range of observational data, from the cosmic microwave background (CMB) and large-scale structure to supernovae luminosity distances and baryon acoustic oscillations [1, 2]. Despite this success, $\Lambda$CDM remains conceptually incomplete and faces several persistent problems. It relies on the presence of dark energy in the form of a cosmological constant ($\Lambda$) and cold dark matter (CDM), two components that dominate the universe’s energy budget but remain elusive in their physical origin [3]. The cosmological constant problem, which involves an enormous discrepancy between the observed value of $\Lambda$ and theoretical predictions from quantum field theory, is a prime example of this tension [4, 5]. Additionally, there are observational inconsistencies such as the Hubble tension—a mismatch between local and early-universe measurements of the Hubble constant [6, 7, 8]—and anomalies in the growth of cosmic structures [9].
@@
The appeal of $f(Q)$ gravity lies in its mathematical simplicity, second-order field equations, and the geometric reinterpretation of gravity. Over the last few years, $f(Q)$ gravity has been investigated across a variety of domains, including early- and late-time cosmology [21], black hole solutions [22], relativistic stars [23], wormholes [24], gravitational wave polarizations [25], and even quantum cosmology [26]. In the cosmological context, $f(Q)$ models can explain cosmic acceleration [27, 28] and alleviate certain observational tensions [29, 28]. In addition, $f(Q)$ allows us to reproduce well-known cosmological scenarios from GR [30] and, from there, to build extensions with new phenomenology and compare them with observations [27]. Another attempt in this direction is the STEGR supplemented with a general power-law term [31], which aimed at accounting for the dark energy. However, this kind of models without cosmological constant is disfavored when confronted with observations [32, 33]. Another example of this is the DGP model with its extensions being constrained by observational data [34]. On the other hand, power-law and exponential forms of $f(Q)$ have been shown to yield accelerated expansion consistent with observations [35, 36].
@@
Several studies have now highlighted the dynamical impact of non-trivial connections in $f(Q)$ cosmology. For instance, in [37] the authors demonstrated that some connection components can become genuinely dynamical, leading to new cosmological solutions beyond GR. The review [38] clarifies further the mathematical underpinnings and implications of the geometrical trinity, emphasizing the promising framework of non-metricity-based theories. Recent works have investigated the appearance of phantom behavior [40], strong coupling and ghost instabilities [44] that can be solved with non-pathological symmetric teleparallel setup [45], or a direct coupling of the matter field to the connection or non-minimal couplings [46], tilt-induced anisotropies [47], and even quantum cosmological corrections via the Wheeler-DeWitt equation [26], all dependent on the chosen connection. Following the above, in [44] and [48], the authors critically assess the viability of $f(Q)$ gravity. Both studies highlight severe theoretical issues, including the presence of ghost degrees of freedom, strong coupling in cosmological perturbations, and the breakdown of standard constraint analysis methods. These pathologies undermine the consistency and predictive power of the theory, especially in early-universe applications. To overcome these limitations, the latter paper proposes extending the framework to $f(R,Q)$ gravity, incorporating both curvature and non-metricity scalars, while [49] identifies a ghost in scalar-nonmetricity theories but shows it to be nonpropagating via ADM analysis and second-class constraints. However, in [44] the authors claim that $f(Q)$ theory propagates the maximum number of degrees of freedom, and this implies there will always be a ghost—not just in cosmological solutions, but in all possible configurations. The remedy comes with [45], where a class of ghost-free models can be constructed in the framework of symmetric teleparallel geometry by enforcing invariance under Transverse Diffeomorphisms (TDiffs), reducing the symmetry group to eliminate problematic modes. Together, these studies offer promising directions to stabilize $f(Q)$-based theories while preserving their geometric appeal and cosmological relevance.
@@
This growing body of work suggests that the affine connection in $f(Q)$ gravity is not a mere auxiliary construct but a physically meaningful component that directly influences observable predictions. It becomes essential, therefore, to systematically study the cosmological consequences of different symmetric and flat connections, especially in homogeneous and isotropic spacetimes. Understanding the role of the connection may help address some of the unresolved issues in $\Lambda$CDM cosmology and provide a framework for constructing more predictive and robust theories of gravity [38].
@@ II Reviewing $f(Q)$ gravitational theory
From now onwards, the units are adopted as $18\pi G=$ 1 and $c=1$. We derive the modified Einstein equation by varying the above action with respect to the metric [63]:
@@ III FLRW Cosmology
The connection for the case (II), denoted by $\Gamma^{\text{(II)}}_{Q}$ reads333 Notice that we have used $\tilde{C}_{2}$ here in order to distinguish it from the rescaled one $C_{2}\equiv\tilde{C}_{2}/a^{2}$. The latter will be used later in the expression of the field equations to follow the notation used in Ref. [37].:
@@
Then, there will be three sets of equations of motion for each of the three possible connections. Let us write explicitly the nonzero terms, and recall that $\mathscr{M}_{\mu\nu}=T_{\mu\nu}$, while [^4] In addition, for each connection, the equations $\mathscr{C}_{r}=0$, $\mathscr{C}_{\theta}=0$ and $\mathscr{C}_{\phi}=0$ are automatically satisfied. $\mathscr{C}_{\alpha}=0$ For the connection $\Gamma^{\text{(I)}}_{Q}$, these are:
@@
which is reduced to $03H(f^{\prime\prime}\dot{Q})+\frac{d}{dt}(f^{\prime\prime}\dot{Q})=$ 0 when we fix $N=1$. Then, we can get:
@@ IV.1.2 Stability of the vacuum solutions analysis
Connection $\Gamma^{\text{(I)}}_{Q}$: The background, i.e. 00th order equations are identically satisfied, while the $1$st order ones will give,
@@
Connection $\Gamma^{\text{(II)}}_{Q}$: As expected, the 00th order equations of the set (28)-(30) are satisfied, while the $1$st order ones give
@@
and the $1$st order equations (namely Eq. (35) - (37)) are satisfied for any $N_{p}(t)$. The non-metricity scalar is unaffected from the perturbations at linear order, i.e. $Q=-6H_{0}^{2}$, ensuring stability.
@@ V.1.3 Connection $\Gamma^{\text{(II)}}_{Q}$
In order to find non-GR solutions, we follow a similar method as we adopted for the connection $\Gamma^{\text{(I)}}_{Q}$. We choose the same initial conditions for $\rho$ and $H$ at $t_{0}=100$ as those in the previous cases. The initial conditions for $Q$ and $\dot{Q}$ are determined by those of $C_{2}$ and $\dot{C}_{2}$ via Eqs. (71) and (72). Then, the initial condition for $\dot{C}_{2}$ is set by assuming that $w_{\textrm{eff}}$ is close to $1/3$ at $t_{0}$. We then numerically integrate Eqs. (27)-(30) backward in $t$. The results are shown in Fig. 3, in which $C_{2}(t_{0})/H(t_{0})$ is set to be $-1/3$, $-1/2$, and $-1$, from blue to cyan. The solutions possess a Big Bang singularity in the past. However, as opposed to the case of connection $\Gamma^{\text{(I)}}_{Q}$ where $\dot{Q}$ diverges at the singularity, the value of $Q$ converges to a non-zero finite value toward the past and its derivative $|\dot{Q}|\rightarrow 0$ there. Therefore, the solutions approach GR-like solutions near the Big Bang singularity in the sense that the effective equation of state approaches $1/3$, and $H^{2}a^{4}$ approaches a non-zero finite value.
@@
Figure 3: The e-folds $N_{e}$ (top) and the ratio $C_{2}/H$ (bottom) as functions of cosmic time $t$ for the cosmology with connection $\Gamma^{\text{(II)}}_{Q}$. Near the Big Bang singularity where $a\rightarrow 0$, we have $H^{2}a^{4}$ approaches a non-zero constant, which implies $H^{2}\propto\rho$. Also, the ratio $C_{2}/H$ scales as $1/a$ near the singularity. In this figure, the initial conditions for $C_{2}(t_{0})/H(t_{0})$ is set to be $-1/3$, $-1/2$, and $-1$, from blue to cyan.
@@ V.2.2 Connection $\Gamma^{\text{(I)}}_{Q}$
To derive the cosmological solutions for connection $\Gamma^{\text{(I)}}_{Q}$, we numerically integrate the field equations (21)-(24) forward in $t$, with initial conditions for $\rho$ and $H$ at $t_{0}=100$ identical to those set in sec. V.2.1. Following a similar method as we did in the case of a radiation-dominated universe, the initial conditions for $Q$ and $\dot{Q}$ are determined by those for $C_{3}$ and $\dot{C}_{3}$ via Eqs. (69) and (70). We then choose the initial conditions for $\dot{C}_{3}$ such that the effective equation of state is close to $w_{\textrm{eff}}\approx-1.2$ at $t=t_{0}$. Finally, the initial conditions for $C_{3}$ are chosen as $H(t_{0})/C_{3}(t_{0})=4$, $3.5$, $3$, $2.5$, and $2$ with results shown in Fig. 5 ($H(t_{0})/C_{3}(t_{0})$ decreases from blue to cyan).
@@
Figure 5: The e-folds $N_{e}(t)$ (top), the ratio $C_{3}/H$ (middle), and the effective equation of state $w_{\textrm{eff}}$ (bottom) as a function of cosmic time $t$ for the phantom-dominated cosmology with connection $\Gamma^{\text{(I)}}_{Q}$. The vertical dashed lines represent the time of the Big Rip singularity. The initial conditions for $C_{3}$ are chosen as $H(t_{0})/C_{3}(t_{0})=4$, $3.5$, $3$, $2.5$, and $2$, from blue to cyan.
@@ V.2.3 Connection $\Gamma^{\text{(II)}}_{Q}$
In Fig. 6, we exhibit the numerical results for the cosmological solutions with the initial conditions $H(t_{0})/C_{2}(t_{0})=5$, $4$, $3$, $2$, and $1$ (from blue to cyan). Similar to the case of $\Gamma^{\text{(I)}}_{Q}$, the non-metricity $Q$ quickly converges to a non-zero constant when $t$ grows. Hence, the solutions approach GR-like and end up with a Big Rip singularity (the dashed vertical lines) where $w_{\textrm{eff}}\rightarrow w_{0}$. In addition, following the same procedure as what we did to obtain Eq. (73), we find that near the singularity, the ratio $C_{2}/H$ converges to
@@
Figure 6: The e-folds $N_{e}(t)$ (top) and the ratio $C_{2}/H$ (bottom) as functions of cosmic time $t$ for the phantom-dominated cosmology with connection $\Gamma^{\text{(II)}}_{Q}$. The vertical dashed lines represent the time of the Big Rip singularity. The initial conditions are chosen as $H(t_{0})/C_{2}(t_{0})=5$, $4$, $3$, $2$, and $1$, from blue to cyan.

### v4

@@ I Introduction
[…]
@@
The appeal of $f(Q)$ gravity lies in its mathematical simplicity, second-order field equations, and the geometric reinterpretation of gravity. Over the last few years, $f(Q)$ gravity has been investigated across a variety of domains, including early- and late-time cosmology [24], inflation [25, 26] , black hole solutions [27], relativistic stars [28], wormholes [29], gravitational wave polarizations [30], and even quantum cosmology [31]. In the cosmological context, $f(Q)$ models can explain cosmic acceleration [32, 33] and alleviate certain observational tensions [34, 33]. In addition, $f(Q)$ allows us to reproduce well-known cosmological scenarios from GR [35] and, from there, to build extensions with new phenomenology and compare them with observations [32]. Another attempt in this direction is the STEGR supplemented with a general power-law term [36], which aimed at accounting for the dark energy. However, such models without cosmological constant are disfavored when confronted with observations [37, 38]. Another example of this is the DGP model with its extensions being constrained by observational data [39]. On the other hand, power-law and exponential forms of $f(Q)$ have been shown to yield accelerated expansion consistent with observations [40, 41].
@@
Several studies have now highlighted the dynamical impact of non-trivial connections in $f(Q)$ cosmology. For instance, in [42] the authors demonstrated that some connection components can become genuinely dynamical, leading to new cosmological solutions beyond GR. The review [43] clarifies further the mathematical underpinnings and implications of the geometrical trinity, emphasizing the promising framework of non-metricity-based theories. Recent works have investigated the appearance of phantom behavior [45], strong coupling and ghost instabilities [49] that can be solved with non-pathological symmetric teleparallel setup [50], or a direct coupling of the matter field to the connection or non-minimal couplings [51], tilt-induced anisotropies [52], and even quantum cosmological corrections via the Wheeler-DeWitt equation [31], all dependent on the chosen connection. Following the above, in [49] and [53], the authors critically assess the viability of $f(Q)$ gravity. Both studies highlight severe theoretical issues, including the presence of ghost degrees of freedom, strong coupling in cosmological perturbations, and the breakdown of standard constraint analysis methods. These pathologies undermine the consistency and predictive power of the theory, especially in early-universe applications. To overcome these limitations, the latter paper proposes extending the framework to $f(R,Q)$ gravity, incorporating both curvature and non-metricity scalars, while [54] identifies a ghost in scalar-nonmetricity theories but shows it to be nonpropagating via ADM analysis and second-class constraints. However, in [49] the authors claim that $f(Q)$ theory propagates the maximum number of degrees of freedom, and this implies there will always be a ghost—not just in cosmological solutions, but in all possible configurations. The remedy comes with [50], where a class of ghost-free models can be constructed in the framework of symmetric teleparallel geometry with the Lagrangian $\mathscr{L}=Q+c\,Q_{\alpha}Q^{\alpha}$. This case is special because the term $Q_{\alpha}Q^{\alpha}$ at least preserves Transverse Diffs (a subgroup of Diffs corresponding to diffeomorphisms with unit Jacobian). Consequently, it is possible to remove ghosts from the spectrum, as in [55], where this class of theories has been analysed.
@@
[…]
@@ II Reviewing $f(Q)$ gravitational theory
From now onwards, the units are adopted as $8\pi G=1$ and $c=1$. We derive the modified Einstein equation by varying the above action with respect to the metric [69]:
@@ III FLRW Cosmology
The connection for the case (II), denoted by $\Gamma^{\text{(II)}}_{Q}$ reads [^3] Notice that we have used $\tilde{C}_{2}$ here in order to distinguish it from the rescaled one $C_{2}\equiv\tilde{C}_{2}/a^{2}$. The latter will be used later in the expression of the field equations to follow the notation used in Ref. [42].:
@@
Then, there will be three sets of equations of motion for each of the three possible connections. Let us write explicitly the nonzero terms, and recall that $\mathscr{M}_{\mu\nu}=T_{\mu\nu}$, while [^4] In addition, for each connection, the equations $\mathscr{C}_{r}=0$, $\mathscr{C}_{\theta}=0$ and $\mathscr{C}_{\phi}=0$ are automatically satisfied. $\mathscr{C}_{\alpha}=0$ . For the connection $\Gamma^{\text{(I)}}_{Q}$, these are:
@@
which is reduced to $3H(f^{\prime\prime}\dot{Q})+\frac{d}{dt}(f^{\prime\prime}\dot{Q})=0$ when we fix $N=1$. Then, we can get:
@@ IV.1.2 Stability of the vacuum solutions analysis
Connection $\Gamma^{\text{(I)}}_{Q}$: The background, i.e. 0th order equations are identically satisfied, while the 11st order ones will give,
@@
Connection $\Gamma^{\text{(II)}}_{Q}$: As expected, the 0th order equations of the set (28)-(30) are satisfied, while the 11st order ones give
@@
and the 11st order equations (namely Eq. (35) - (37)) are satisfied for any $N_{p}(t)$. The non-metricity scalar is unaffected from the perturbations at linear order, i.e. $Q=-6H_{0}^{2}$, ensuring stability.
@@ V.1.3 Connection $\Gamma^{\text{(II)}}_{Q}$
In order to find non-GR solutions, we follow a similar method as we adopted for the connection $\Gamma^{\text{(I)}}_{Q}$. We choose the same initial conditions for $\rho$ and $H$ at $t_{0}=100$ as those in the previous cases. The initial conditions for $Q$ and $\dot{Q}$ are determined by those of $C_{2}$ and $\dot{C}_{2}$ via Eqs. (71) and (72). Then, the initial condition for $\dot{C}_{2}$ is set by assuming that $w_{\textrm{eff}}$ is close to $1/3$ at $t_{0}$. We then numerically integrate Eqs. (27)-(30) backward in $t$. The results are shown in Fig. 3, in which $C_{2}(t_{0})/H(t_{0})$ is set to be 1/3-1/3, 1/2-1/2, and 1-1, from blue to cyan. The solutions possess a Big Bang singularity in the past. However, as opposed to the case of connection $\Gamma^{\text{(I)}}_{Q}$ where $\dot{Q}$ diverges at the singularity, the value of $Q$ converges to a non-zero finite value toward the past and its derivative $|\dot{Q}|\rightarrow 0$ there. Therefore, the solutions approach GR-like solutions near the Big Bang singularity in the sense that the effective equation of state approaches $1/3$, and $H^{2}a^{4}$ approaches a non-zero finite value.
@@
Figure 3: The e-folds $N_{e}$ (top) and the ratio $C_{2}/H$ (bottom) as functions of cosmic time $t$ for the cosmology with connection $\Gamma^{\text{(II)}}_{Q}$. Near the Big Bang singularity where $a\rightarrow 0$, we have $H^{2}a^{4}$ approaches a non-zero constant, which implies $H^{2}\propto\rho$. Also, the ratio $C_{2}/H$ scales as $1/a$ near the singularity. In this figure, the initial conditions for $C_{2}(t_{0})/H(t_{0})$ is set to be 1/3-1/3, 1/2-1/2, and 1-1, from blue to cyan.
@@ V.2.2 Connection $\Gamma^{\text{(I)}}_{Q}$
To derive the cosmological solutions for connection $\Gamma^{\text{(I)}}_{Q}$, we numerically integrate the field equations (21)-(24) forward in $t$, with initial conditions for $\rho$ and $H$ at $t_{0}=100$ identical to those set in sec. V.2.1. Following a similar method as we did in the case of a radiation-dominated universe, the initial conditions for $Q$ and $\dot{Q}$ are determined by those for $C_{3}$ and $\dot{C}_{3}$ via Eqs. (69) and (70). We then choose the initial conditions for $\dot{C}_{3}$ such that the effective equation of state is close to $w_{\textrm{eff}}\approx-1.2$ at $t=t_{0}$. Finally, the initial conditions for $C_{3}$ are chosen as $H(t_{0})/C_{3}(t_{0})=4$, $3.5$, 33, $2.5$, and $2$ with results shown in Fig. 5 ($H(t_{0})/C_{3}(t_{0})$ decreases from blue to cyan).
@@
Figure 5: The e-folds $N_{e}(t)$ (top), the ratio $C_{3}/H$ (middle), and the effective equation of state $w_{\textrm{eff}}$ (bottom) as a function of cosmic time $t$ for the phantom-dominated cosmology with connection $\Gamma^{\text{(I)}}_{Q}$. The vertical dashed lines represent the time of the Big Rip singularity. The initial conditions for $C_{3}$ are chosen as $H(t_{0})/C_{3}(t_{0})=4$, $3.5$, 33, $2.5$, and $2$, from blue to cyan.
@@ V.2.3 Connection $\Gamma^{\text{(II)}}_{Q}$
In Fig. 6, we exhibit the numerical results for the cosmological solutions with the initial conditions $H(t_{0})/C_{2}(t_{0})=5$, $4$, 33, 22, and $1$ (from blue to cyan). Similar to the case of $\Gamma^{\text{(I)}}_{Q}$, the non-metricity $Q$ quickly converges to a non-zero constant when $t$ grows. Hence, the solutions approach GR-like and end up with a Big Rip singularity (the dashed vertical lines) where $w_{\textrm{eff}}\rightarrow w_{0}$. In addition, following the same procedure as what we did to obtain Eq. (73), we find that near the singularity, the ratio $C_{2}/H$ converges to
@@
Figure 6: The e-folds $N_{e}(t)$ (top) and the ratio $C_{2}/H$ (bottom) as functions of cosmic time $t$ for the phantom-dominated cosmology with connection $\Gamma^{\text{(II)}}_{Q}$. The vertical dashed lines represent the time of the Big Rip singularity. The initial conditions are chosen as $H(t_{0})/C_{2}(t_{0})=5$, $4$, 33, 22, and $1$, from blue to cyan.
@@
Before closing this section, we would like to mention that the existence of the Big Bang and Big Rip singularities in the cosmological solutions we found for the connections $\Gamma^{\text{(I)}}_{Q}$ and $\Gamma^{\text{(II)}}_{Q}$ are quite robust against different choices of initial conditions at $t_{0}$, as long as the cosmological behaviors at $t_{0}$ are close to that in standard GR cosmology. However, we still cannot completely exclude the possibility that there may be some fine-tuned initial conditions such that the singularities are resolved for these connections. A thorough analysis of the dependence between the choices of initial conditions and the resulting cosmological solutions can be carried out efficiently via dynamical analysis. We will leave this for future work.